### v2

# Counterspeech the ultimate shield! Multi-Conditioned Counterspeech Generation through Attributed Prefix Learning

Abstract
Counterspeech has proven to be a powerful tool to combat hate speech online. Previous studies have focused on generating counterspeech conditioned only on specific intents  [^1] While ‘intent’ has an established meaning in Natural Language Generation, we retain this term to maintain consistency with the IntentCONANv2. (single attributed). However, a holistic approach considering multiple attributes simultaneously can yield more nuanced and effective responses. Here, we introduce HiPPrO, Hierarchical Prefix learning with Preference Optimization, a novel two-stage framework that utilizes the effectiveness of attribute-specific prefix embedding spaces hierarchically optimized during the counterspeech generation process in the first phase. Thereafter, we incorporate both reference and reward-free preference optimization to generate more constructive counterspeech. Furthermore, we extend IntentCONANv2 by annotating all $13,973$ counterspeech instances with emotion labels by five annotators. HiPPrO leverages hierarchical prefix optimization to integrate these dual attributes effectively. An extensive evaluation demonstrates that HiPPrO achieves a $\sim 38\%$ improvement in intent conformity and a $\sim 3\%$, $\sim 2\%$, $\sim 3\%$ improvement in Rouge-$1$, Rouge-$2$, and Rouge-L, respectively, compared to several baseline models. Human evaluations further substantiate the superiority of our approach, highlighting the enhanced relevance and appropriateness of the generated counterspeech. This work underscores the potential of multi-attribute conditioning in advancing the efficacy of counterspeech generation systems.222 Warning: The materials presented in this paper might be disturbing or offensive.  [^3] Accepted in ACL 2025 Main Conference.
Counterspeech the ultimate shield! Multi-Conditioned Counterspeech Generation through Attributed Prefix Learning
Aswini Kumar1, Anil Bandhakavi2, Tanmoy Chakraborty1 1IIT Delhi, India, 2Logically.ai { aswinikumarpadhi1995@gmail.com, } anil@logically.ai, tanchak@iitd.ac.in
1 Introduction
Figure 1: An illustration of the output of existing methods in generating effective counterspeech, HiPPrO (ours) demonstrates superior performance in producing high-quality, effective, multi-attributed counterspeech for a given hate speech without relying on instructional prompts.
The rise of the Internet has transformed social media platforms into hostile grounds for hateful comments targeting specific communities. Hate speech (HS) (Awal et al., 2021; Chakraborty and Masud, 2022) carries offensive statements that leverage stereotypes to spread harmful content. In contrast, counterspeech (CS) Benesch et al. (2016a); Wright et al. (2017) involves constructive responses to counteract hate speech by promoting positive dialogue, thus mitigating online hostility while preserving diverse perspectives Schieb and Preuss (2016); Chandrasekharan et al. (2017). As hateful comments surge Leetaru (2019); Masud et al. (2021, 2022), depending only on human-generated counterspeech becomes impractical. In this context, machine learning models appear essential for analyzing and generating counterspeech, offering a promising solution for automating the process. By leveraging this, content moderators can efficiently counter online hate Parker and Ruths (2023); Garg et al. (2023); Chung et al. (2023). Consequently, several interests Mathew et al. (2019); Qian et al. (2019); Fanton et al. (2021a); Bonaldi et al. (2022); Hengle et al. (2024) have intensified in the development of counterspeech generation methods.
The conceptualization of CS generation has evolved from a simple sequence-to-sequence problem to a more nuanced approach, acknowledging the diverse and context-dependent nature of hate speech Chung et al. (2019); Mathew et al. (2019); Sheng et al. (2020); Parker and Ruths (2023); Chung et al. (2023). This paradigm shift has fostered the development of advanced generative models incorporating stylistic and condition-guided elements, such as politeness, joyfulness, and detoxification, to create more effective counter-narratives Saha et al. (2022); Sheng et al. (2020). Recent research has introduced intent-specific CS generation, where established strategies guide the generation process to combat hate speech Gupta et al. (2023); Hengle et al. (2024); Benesch et al. (2016b).
Motivation:
A significant portion of online hate speech consists of short, abusive statements Benesch et al. (2016a), and CS has shown potential in effectively countering such harmful content. While conventional approaches generate a single response per hate speech instance Zhu and Bhat (2021); Qian et al. (2019), recent studies highlight the benefits of tailoring CS to specific attributes for generating more diverse responses Gupta et al. (2023); Hengle et al. (2024). In practical scenarios, hateful comments often include multiple user intentions, necessitating the development of counterspeech that effectively addresses diverse attributes, resulting in more comprehensive and effective responses. For instance, single-attribute approaches typically produce one-dimensional responses by focusing solely on intent (e.g., being informative) while neglecting the emotional harmony required for persuasive communication. Real-world hate speech frequently needs a range of emotional responses, making it imperative to generate counterspeech that balances factual accuracy with emotional engagement.
The integration of Large Language Models (LLMs) for various text generation tasks has become increasingly popular Yang et al. (2024), but training these models is resource intensive. To address this, parameter-efficient fine-tuning (PEFT) techniques, such as tunable prefixes Li and Liang (2021), have gained popularity. These techniques involve adding task-specific continuous vectors (key-value pairs) to transformer layers while keeping the rest of the model unchanged. Recent studies Liu et al. (2023a) showed that these vectors excel in generating conditional text by capturing hidden implied relationships during training. In this work, we propose generating multi-attribute guided counterspeech generation through a hierarchical approach to learning continuous prefix vectors, enabling more varied and contextually relevant responses. We provide an example in Figure 1, where our method, HiPPrO, is compared to leading models like Generate-Prune-Select (GPS) Zhu and Bhat (2021), DialoGPT Zhang et al. (2020b), and COARL Hengle et al. (2024). While traditional models produce semantically sound responses, they often lack nuance and struggle with complex contextual relationships, as mentioned in (Benesch et al., 2016a). Inspired by Liu et al. (2023b), HiPPrO enables to generate more effective and contextually relevant counterspeech by analyzing hate speech and multiple user intentions, resulting in more impactful and persuasive responses.
Our Contribution:
This study introduces an advanced pipeline for counterspeech generation, addressing the implicit nature of hate speech with responses aligned to multiple attributes. We focus on four primary intents – ‘positive’, ‘informative’, ‘questioning’, and ‘denouncing.’ and five emotion categories –‘anger’, ‘disgust’, ‘joy’, ‘sadness’, and ‘surprise’. Furthermore, we curate MultiCONAN, the largest collection of intent-emo-specific counterspeech, with $13,952$ responses countering $3,487$ instances of hate speech. We propose HiPPrO, a novel two-phase framework that first learns attribute-specific prefix embeddings (key-value pairs) and then applies preference tuning to generate constructive, non-toxic responses. Comprehensive evaluations using automated metrics and human evaluations demonstrate that HiPPrO consistently outperforms existing methods in CS generation across various criteria and achieves comparable performance with state-of-the-art LLMs like GPT-3.5 and GPT-4.444The source codes and datasets attached as supplementary and will be made public upon acceptance of the paper.
2 Related Work
The evolution of CS datasets has progressed from crowdsourced collections to expert-curated, intent-specific compilations Qian et al. (2019); Chung et al. (2019); Fanton et al. (2021a); Gupta et al. (2023); Hengle et al. (2024), while CS generation techniques have advanced from basic sequence-to-sequence models to sophisticated multi-phase pipelines with attribute control Zhu and Bhat (2021); Saha et al. (2019). These advancements have significantly improved the nuance and effectiveness of automated CS in promoting constructive dialogue Benesch et al. (2016c). Concurrently, parameter-efficient fine-tuning methods like Prefix Tuning Li and Liang (2021) and prompt tuning Lester et al. (2021) have emerged, modifying inputs while preserving language model parameters. Recent developments in preference tuning include RLHF’s application to instruction-following tasks Ouyang et al. (2022a), Direct Preference Optimization (DPO) Rafailov et al. (2023), and a unified approach omitting both reward and reference models Hong et al. (2024), addressing challenges in scalability and model sensitivity.
Counterspeech Intent and Emotion
Target
IN
and
AN
IN
and
DI
IN
and
JO
IN
and
SA
IN
and
SU
DE
and
AN
DE
and
DI
DE
and
JO
DE
and
SA
DE
and
SU
PO
and
AN
PO
and
DI
PO
and
JO
PO
and
SA
PO
and
SU
QU
and
AN
QU
and
DI
QU
and
JO
QU
and
SA
QU
and
SU
Total
Muslim 236 269 248 149 19 221 420 232 33 9 187 55 622 54 1 397 147 24 9 337 3669
Women 108 159 161 71 10 151 181 152 14 10 59 44 378 26 1 258 91 15 11 134 2034
LGBT+ 93 171 129 48 8 140 208 85 12 4 45 36 335 31 2 190 96 11 15 137 1796
Jews 123 113 94 53 10 109 206 54 14 9 73 32 251 36 1 151 67 5 5 164 1570
Refugee 10 8 37 12 3 10 22 33 5 0 5 4 55 6 0 27 11 3 1 28 280
AP 7 10 6 5 1 9 16 4 0 0 4 1 22 2 0 11 4 1 2 11 116
IMGT 99 101 260 87 15 147 137 244 29 5 64 22 441 32 3 242 76 30 14 200 2248
Disable 30 51 65 28 1 33 75 59 7 1 12 17 125 18 1 100 18 6 7 42 696
PoC 62 122 68 47 3 80 148 54 13 7 45 13 223 20 1 122 67 8 5 100 1208
Others 25 23 28 11 2 28 37 16 6 2 19 7 57 5 1 38 19 3 1 28 356
Total 793 1027 1096 511 72 928 1450 933 133 47 513 231 2509 230 11 1536 596 106 70 1181 13973
Train 505 656 686 325 39 605 912 544 84 30 294 142 1594 132 7 978 381 68 47 772 8801
Dev 122 158 161 90 10 152 247 140 29 6 75 32 425 49 3 235 87 11 10 159 2201
Test 166 213 249 96 23 171 291 249 20 11 144 57 490 49 1 323 128 27 13 250 2971
Table 1: Counterspeech distribution in MultiCONAN across various multi-attribute combinations, categorized by target groups (see abbreviations in Section 3).
3 The MultiCONAN Dataset
Due to the superior quality of counterspeech (see Appendix 8.3), we selected IntentCONANv2 Hengle et al. (2024) as the foundation for our work. IntentCONANv2, while valuable, has limitations that restrict its ability to fully capture the diversity and complexity of counterspeech. Its dependence on a single attribute, ‘intent,’ oversimplifies responses and fails to account for the emotional tone, which plays a critical role in shaping effective counterspeech. For instance, an informative counterspeech can convey vastly different emotional tones, such as joy or sadness, which are not captured in the original dataset. Building upon this, we introduce MultiCONAN, an enhanced version featuring additional emotion class labels for each counterspeech instance. MultiCONAN includes $13,973$ CS instances of IntentCONANv2, each tagged with one of five emotion classes: Anger (AN), Joy (JO), Disgust (DI), Sad (SA), and Surprise (SU) (See Table 1). The added emotion annotations enhance analysis granularity and facilitate the development of models that integrate both intent and emotional context. This annotation framework enables exploration of how emotional tone interacts with strategic intent and supports the creation of nuanced CS generation models Gupta et al. (2023); Hengle et al. (2024). MultiCONAN thus serves as a valuable resource for advancing research in CS generation, aiming to produce contextually appropriate and emotionally resonant responses. For detailed information on the annotation process and the statistics of the data set, refer to Appendices 8.1,8.2, and 8.4, respectively.
Figure 2: Our proposed model, HiPPrO, follows a two-phase pipeline. In Phase 1, we train the prefix parameters ($\alpha,\beta,\gamma,\delta$) associated with prefix adapters ($\mathcal{F_{\alpha}},\mathcal{F_{\beta}}$) on encoder side and ($\mathcal{F_{\gamma}},\mathcal{F_{\delta}}$) on the decoder side. For $\alpha$ and $\beta$, the model is trained with hate speech and intent, separated by </s> as input. While training $\gamma$ and $\delta$, the optimal parameters, $\alpha^{*}$ and $\beta^{*}$, are kept fixed, and it includes hate speech, intent, and emotion, separated by </s> as input $x$, this is also used as a prompt input during the second phase. In Phase 2, we apply preference tuning using an odd ratio loss, where the ground-truth counterspeech serves as the chosen candidate $y_{s}$, and the model-generated counterspeech is treated as the rejected candidate $y_{r}$.
4 Proposed Methodology
In this section, we elaborate on the inner workings and structural components of HiPPrO, a novel automated counterspeech generation framework. Here, we explain how our designed model can address the previous challenges by (i) generating multi-intentional conditioned counterspeech that can address hateful comments through semantic relevancy, and (ii) aligning it with qualitative human-generated responses through a reward and reference-model-free approach to ensure that the generation is contrastive (Figure 2).
Task Formulation
We use our curated MultiCONAN dataset for generating multi-intentional counterspeech generation. Considering our dataset as $\mathcal{D}=\{(h_{1},i_{1},e_{1},c_{1}),\ldots,(h_{n},i_{n},e_{n},c_{n})\}$, where $h_{i}\in\mathcal{H}$ is the $i$-th hate speech statement, $c_{i}\in\mathcal{C}$ is the counterspeech corresponding to $h_{i}$, and $i_{i}\in\mathcal{I}$ and $e_{i}\in\mathcal{E}$ are the intent and emotion categories of $c_{i}$, respectively. Our objective is to learn a stochastic counterspeech generation function $\psi:\mathcal{H}\times\mathcal{I}\times\mathcal{E}\rightarrow\mathcal{C}$, such that $c_{i}\sim\psi(\cdot|h_{i},i_{i},e_{i})$.
We address this problem by decomposing the counterspeech generation task into two phases. In the first phase, we focus on learning the prefix vectors for individual attributes and capturing their conditional dependencies using a two-step method. Initially, we learn the prefix vector for the intent category and collect the optimal prefix vectors. Subsequently, we add another prefix vector initialized with the previously learned values and optimized it for both the intent and emotion categories while keeping the model parameters from step one fixed. In the second phase, we employ a reward and reference model-free alignment method called the odds ratio preference optimization algorithm Hong et al. (2024). During this, we treat the counterspeech generated in phase one as the rejected data column and the actual ground truth counterspeech as the selected data column. For all our experiments, we utilize FLAN-T5 Chung et al. (2022) as the base model due to its robust reasoning and multi-task learning capabilities.
Phase 1: Hierarchical Prefix Optimization (HIPO)
In this section, we mathematically formalize the concept of hierarchical prefix learning. The prefix vectors, represented by tunable key-value pairs Vaswani et al. (2017), are introduced in two sub-phases. In the first sub-phase, we add $|V_{\mathcal{I}}|$ virtual prefixes with dimension $d$ across $l$ layers. The prefix adapters, $\mathcal{F_{\alpha}}$ and $\mathcal{F_{\beta}}$, introduce task-specific continuous vectors, $\alpha,\beta\in\mathbb{R}^{|V_{\mathcal{I}}|\times l\times 2d}$ to the encoder and decoder, respectively, guiding counterspeech generation according to user intent.
We maximize the expected log-likelihood and collect the optimal prefix adapters $\mathcal{F_{\alpha^{*}}},\mathcal{F_{\beta^{*}}}$ for the intent-guided counterspeech generation, where the hate speech and intent are sampled from $\mathcal{D}$, as denoted by,
$\mathcal{F_{\alpha^{*}}},\mathcal{F_{\beta^{*}}}=\underset{\alpha,\beta}{\mathrm{argmax}}\mathop{{}\mathbb{E}_{(h,i\sim\mathcal{D})}}\log(\mathcal{F_{\alpha}}[\pi^{ENC}_{\theta}(X(h,i);$ (1)
$\theta,\alpha)],\mathcal{F_{\beta}}[\pi^{DEC}_{\theta}(X(c);\theta,\beta)])$
where $\pi^{ENC}_{\theta}$ and $\pi^{DEC}_{\theta}$ are the encoder and decoder part of the model, respectively. The input $X=[h;i;c]$, where $(h,i,c)\in\mathcal{D}$. We further add another set of adapters, $\mathcal{F_{\gamma}}$, and $\mathcal{F_{\delta}}$, on the encoder and decoder side, which add trainable prefix vectors $\gamma,\delta\in\mathbb{R}^{|V_{\mathcal{E}}|\times l\times 2d}$ on top of the previously trained $\mathcal{F_{\alpha^{*}}},\mathcal{F_{\beta^{*}}}$, where $|V_{\mathcal{E}}|$ is the number of virtual tokens for second sub-phase. These prefix vectors are responsible for guiding the counterspeech generation with intent and emotion-specific attributes. Here, the reformulated input is $X^{{}^{\prime}}=[h;i;e;c]$, where $(h,i,e,c)\in\mathcal{D}$. During training, we freeze the model parameters $\theta$ and the intent-specific prefix parameters $\alpha,\beta$ and maximize the expected log-likelihood to get optimal prefix adapters $\mathcal{F_{\gamma^{*}}},\mathcal{F_{\delta^{*}}}$ for optimal $\gamma^{*},\delta^{*}$ given by,
$(\displaystyle\mathcal{F_{\gamma^{*}}},\mathcal{F_{\delta^{*}}}=\underset{\gamma,\delta}{\mathrm{argmax}}\mathop{{}\mathbb{E}_{(h,i,e\sim\mathcal{D})}}\log(\mathcal{F_{\gamma}}[\mathcal{F_{\alpha^{*}}}[\pi^{ENC}_{\theta}($2)
$X^{{}^{\prime}}(h,i,e);\theta,\alpha^{*},\gamma)]],\mathcal{F_{\delta}}[\mathcal{F_{\beta^{*}}}[\pi^{DEC}_{\theta}(X^{{}^{\prime}}(c);$
$\theta,\beta^{*},\delta)]])$
Here, all $\alpha^{*},\beta^{*},\gamma^{*}$ and $\delta^{*}$ are responsible for the intent and emotion-conditioned counterspeech generation. Now, our model is ready to advance to Phase 2, where we apply a preference-tuning approach to all pre-trained prefix vectors. This step further optimizes the model to align counterspeech responses more closely with human-generated outputs.
Phase 2: Preference Optimization
In this phase, we focus on optimizing our model to produce counterspeech instances that are both effective and non-toxic by employing a reward and reference-free preference tuning approach. Inspired by Hong et al. (2024), we use the odd ratio method to align the output with the ground truth preferences. This method self-penalizes the model output probabilities using its corresponding odds denoted as,
$Odds_{\alpha,\beta,\gamma,\delta}(y|x)=\frac{\pi_{\theta,\alpha,\beta,\gamma,\delta}(y|x)}{1-\pi_{\theta,\alpha,\beta,\gamma,\delta}(y|x)}$ (3)
Now this ratio has an interesting characteristic – it boils down to less than one when the desired probability is less than its odds, which eventually penalizes the model output and gives a smaller number. The odd ratio is the ratio between two odds of two independent events. The odd ration $OR$ is given by,
$OR_{\alpha,\beta,\gamma,\delta}(y_{s},y_{r})=\frac{Odds_{\alpha,\beta,\gamma,\delta}(y_{s}|x)}{Odds_{\alpha,\beta,\gamma,\delta}(y_{r}|x)}$ (4)
Here, $y_{s}$ denotes the ground-truth counterspeech, which is human-generated, and $y_{r}$ denotes the model-generated counterspeech, which is trained during the first stage. $OR_{\alpha,\beta,\gamma,\delta}(y_{s},y_{r})$ becomes high when the numerator is higher than the denominator, and hence $\pi_{\theta,\alpha,\beta,\gamma,\delta}(y_{s}|x)$.
Let us consider the preference tuning dataset $\mathcal{D^{{}^{\prime}}}=\{(h_{1},y_{s1},y_{r1}),\ldots,(h_{n},y_{sn},y_{rn})\}$, where $y_{si}$ and $y_{ri}$ are the $i^{th}$ ground-truth response and model-generated response for a hate speech $h_{i}$. The final loss is the expectation of prefix-tuned loss and odds ratio loss over the data samples sampled from $\mathcal{D^{{}^{\prime}}}$. As shown in Equation 5, the final loss for preference optimization $J_{final}$ is the combined loss for finetuned loss ($J_{finetuned}$) and odds ratio loss $J_{OR}$ weighted by a factor $\epsilon$.
$J_{final}=\mathop{{}\mathbb{E}_{(x,y_{s},y_{r}\sim\mathcal{D^{{}^{\prime}}})}}[J_{finetuned}+\epsilon J_{OR}]$ (5)
The $J_{finetuned}$ is simply the negative log-likelihood loss obtained in the prefix-tuning step, and $J_{OR}$ is given by,
$J_{OR}=-\log(\sigma(\log(\frac{Odds_{\alpha,\beta,\gamma,\delta}(y_{s}|x)}{Odds_{\alpha,\beta,\gamma,\delta}(y_{r}|x)})))$ (6)
where $\sigma$ denotes the sigmoid function. The final objective is to minimize $J_{final}$ by updating the continuous prefix vectors so that the model outputs can align more with the ground-truth counterspeech. For detailed information regarding computational resources and hyperparameter settings, please refer to Appendix 8.8 and Appendix 8.9, respectively. 555 The model was trained for up to 50 epochs using a callback, with a fixed training batch size and a learning rate of 4 and $1\times e^{-4}$, respectively.
Method Prompt/Adapter ROUGE (↑) M (↑) BS (↑) CoSim (↑) IC (↑) EC (↑) TC (↑) T (↓)
R1 R2 RL
GPS -- $0.089$ $0.011$ $0.075$ $0.055$ $0.840$ $0.287$ $0.247$ $0.292$ $0.394$ $0.146$
DialoGPT -- $0.164$ $0.050$ $0.108$ $0.198$ $0.749$ $0.536$ $0.714$ $0.571$ $0.969$ $0.309$
CoARL -- $0.156$ $0.035$ $0.126$ $0.110$ $0.860$ $0.440$ $0.380$ $0.468$ $0.938$ $0.258$
Vanilla FLAN-T5XXL ZS $0.176$ $0.045$ $0.145$ $0.118$ $0.866$ $0.479$ $0.391$ $0.586$ $0.989$ $0.341$
Vanilla FLAN-T5XXL FS $0.174$ $0.042$ $0.142$ $0.116$ $0.864$ $0.456$ $0.352$ $0.478$ $0.976$ $0.301$
GPT-3.5-Turbo ZS $0.239$ $0.052$ $0.160$ $0.249$ $0.868$ $0.585$ $0.533$ $0.783$ $0.978$ $0.0007$
GPT-3.5-Turbo FS $0.242$ $0.059$ $0.161$ $0.249$ $0.868$ $0.564$ $0.309$ $0.527$ $0.888$ $0.012$
GPT-4 ZS $0.221$ $0.037$ $0.145$ $0.229$ $0.864$ $0.583$ $0.545$ $0.790$ $0.976$ $0.030$
GPT-4 FS $0.226$ $0.039$ $0.149$ $0.213$ $0.868$ $0.554$ $0.454$ $0.515$ $0.932$ $0.014$
Vanilla FLAN-T5 XXL PrefixTuning $0.229$ $0.052$ $0.158$ $0.222$ $0.870$ $0.539$ $0.470$ $0.666$ $0.901$ $0.024$
Vanilla BART Large PrefixTuning $0.207$ $0.042$ $0.131$ $0.226$ $0.861$ $0.226$ $0.249$ $0.453$ $0.915$ $0.030$
Vanilla GPT2 XL PrefixTuning $0.155$ $0.029$ $0.122$ $0.084$ $0.820$ $0.487$ $0.453$ $0.455$ $0.937$ $0.837$
Vanilla Llama 3.1 Instruct 8B PrefixTuning $0.168$ $0.043$ $0.135$ $0.093$ $0.854$ $0.515$ $0.552$ $0.592$ $0.903$ $0.605$
Vanilla Mistral Instruct 7B PrefixTuning $0.170$ $0.046$ $0.140$ $0.117$ $0.858$ $0.510$ $0.516$ $0.597$ $0.920$ $0.645$
HiPPrO ${}_{\mathrm{\textbf{VT=3}}}$ (Ours) PrefixTuning $0.273^{*}$ $0.082^{*}$ $0.199^{*}$ $0.242$ $0.879^{*}$ $0.567$ $0.929^{*}$ $0.706$ $0.897$ $0.087$
- With out ORPO PrifixTuning $0.272$ $0.081$ $0.198$ $0.241$ $0.879$ $0.567$ $0.928$ $0.705$ $0.896$ $0.111$
- With DPO PrefixTuning $0.272$ $0.081$ $0.198$ $0.240$ $0.879$ $0.566$ $0.928$ $0.685$ $0.895$ $0.089$
- HIPO VT=5 PrifixTuning $0.275$ $0.081$ $0.200$ $0.241$ $0.880$ $0.578$ $0.921$ $0.656$ $0.916$ $0.096$
- HIPO VT=7 PrefixTuning $0.273$ $0.084$ $0.199$ $0.24$ $0.879$ $0.573$ $0.937$ $0.643$ $0.888$ $0.084$
- HIPO VT=10 PrefixTuning $0.271$ $0.081$ $0.197$ $0.244$ $0.878$ $0.564$ $0.905$ $0.630$ $0.858$ $0.077$
$\Delta_{\mathrm{HiPPrO(Ours)}-\mathrm{BestBaselineMethod}}$ $\uparrow 0.034$ $\uparrow 0.023$ $\uparrow 0.038$ $\downarrow 0.007$ $\uparrow 0.009$ $\downarrow 0.018$ $\uparrow 0.384$ $\downarrow 0.084$ $\downarrow 0.092$ $\downarrow 0.086$
Table 2: Comparing HiPPrO with baselines across various evaluation metrics. Here, ↑ (resp. ↓) denotes that higher (resp. lower) is better. Bold (resp. underline) indicates the best (resp. second-ranked) performance. * shows our model significantly outperforms ($p<0.05$) the best baselines – GPT-3.5-Turbo ZS and FS (see Appendix 8.5).
5 Experimental Setup
5.1 Baselines
To evaluate the efficacy of various models, we investigate Generate Prune Select (GPS) Zhu and Bhat (2021), a three-stage pipeline encompassing autoencoding, grammatical filtering, and response selection. Furthermore, we optimize DialoGPT Zhang et al. (2020b) to produce contextually coherent responses. Additionally, we incorporate CoARL Hengle et al. (2024), the state-of-the-art intent-conditioned CS generation approach. Furthermore, we explore prefix tuning on Vanilla FLAN-T5XXL, and BARTLarge (excluding HiPPrO) and conduct experiments with HIPO (without preference optimization) using different virtual token (VT) sizes (VT = $3,5,7,10$), VT = $3$ emerging as the optimal configuration in terms of both parameter efficiency ($589,824$ trainable parameters, $0.0052\%$ of total model parameters) and performance metrics. We consider VT = $3$ for preference tuning. With same setup, we experiment with decoder-only models like GPT2XL, Llama 3.1 Instruct8B and Mixtral Instruct7B. We also utilize DPO to enhance our evaluation further. Our comprehensive assessment encompasses both zero-shot and few-shot performances on three LLMs: Vanilla FLAN-T5XXL Chung et al. (2022), GPT-3.5-Turbo (ChatGPT) and GPT-4 Ouyang et al. (2022b) (see Appendix 8.6 and 8.7). In all the above setups we do not consider natural language instruction prompts.
5.2 Evaluation Metrics
Evaluating CS generation presents challenges due to its dynamic nature, diverse response possibilities, and absence of standardized metrics Chung et al. (2023), prompting our framework to utilize comprehensive, multi-dimensional evaluation metrics. These evaluation metrics include lexical similarity, semantic similarity or relevance, intent conformity, emotion conformity, target conformity, and toxicity score. Lexical similarity is evaluated using Rouge Lin (2004) and Meteor (M) Banerjee and Lavie (2005), which quantify the linguistic alignment between generated and reference texts. Semantic relevance is assessed with cosine similarity (CoSim) Reimers and Gurevych (2019) and BERTScore (BS) Zhang et al. (2020a), ensuring that generated CS engages meaningfully with the primary topic of hate speech. A low relevance score implies a lack of topical coherence, where the CS fails to adequately address the primary subject of hate speech. We also evaluate the effectiveness of incorporating intention, emotional resonance, and target alignment through Intent Conformity (IC) Gupta et al. (2023), Emotion Conformity (EC), and Target Conformity (TC), respectively. These metrics are particularly valuable in scenarios where ground-truth counterspeech is unavailable for unseen hate speech instances. We first train three distinct RoBERTa-large models on our dataset to measure the IC, EC, and TC scores. The models achieve testing accuracy of $0.86$, $0.75$, and $0.88$, respectively, before being considered as evaluation metrics. Toxicity (T) [^6] https://www.perspectiveapi.com/ levels of generated CS using the Hanu and Unitary team (2020) library, ensure that our approach promotes respectful and safe communication.
6 Experimental Results
This section presents a comprehensive empirical analysis that systematically evaluates the efficacy of HiPPrO in comparison to existing state-of-the-art techniques.
6.1 Quantitative Results
Table 2 demonstrates the quantitative evaluation across various metrics. HiPPrO shows a notable improvement over baselines across several evaluation metrics. HiPPrO achieves $0.273$ ROUGE-1, $0.082$ ROUGE-2, $0.199$ ROUGE-L, substantially higher than GPS, DialoGPT, and CoARL, with an average improvement of $0.117$ in ROUGE-1, $0.05$ in ROUGE-2 and $0.096$ in ROUGE-L. This indicates that
HiPPrO’s counterspeech better aligns with reference content in terms of coverage and detail. In BERTScore, HiPPrO scores $0.879$, surpassing GPS ($0.840$), DialoGPT ($0.749$), and CoARL ($0.860$), GPT-4 Few-Shot (FS) ($0.868$) and GPT-3.5 Few-Shot ($0.868$), which indicates HiPPrO’s proficiency in preserving the underlying meaning of the original CS. HiPPrO exhibits a slight decrease in CoSim compared to models like GPT-3.5-Turbo Zero-Shot (ZS) with a score of $0.585$, indicating a potential trade-off between contextual relevance and semantic alignment with reference responses.
The generated CS should ideally be a balanced generation of both intent and emotional attributes, ensuring that responses not only address the harmful content effectively but also align with intent and emotional tone along with the target audience. However, the evaluation results indicate that it is a challenge for many models to achieve this balance. For instance, while GPT-3.5-Turbo ZS demonstrates a low IC score of $0.533$, its EC score is significantly high at $0.783$, highlighting the difficulty in generating CS that is both intent-aligned and emotionally resonant. Similarly, GPT-4 ZS shows a high EC score of $0.790$, yet its IC score of $0.545$ suggests that even more advanced LLMs struggle to maintain emotional alignment in the zero-shot and few-shot scenarios. In contrast, HiPPrO demonstrates a better balance, achieving an IC score of $0.929$ and an EC score of $0.706$. HiPPrO’s IC score is substantially higher, indicating its superior ability to generate intent-aligned CS while still maintaining a better level of emotional resonance. This suggests that HiPPrO is more adept at maintaining the stability between these two critical attributes during the generation process.
The relationship between TC and BS provides insights into how models generate counterspeech; a high TC score with a relatively low BS indicates that the CS is target-specific but may lack deeper semantic alignment with the original CS. For instance, GPT-3.5-Turbo in the ZS setting achieves a TC score of $0.978$ and a BS score of $0.868$, suggesting that while it effectively mentions the target group, it may produce overly generic responses that fail to address nuanced aspects of hate speech. In contrast, HiPPrO manages to strike a better balance by achieving a TC score of $0.897$, which, while slightly lower than GPT-4 few-shot, is complemented by a BS of $0.879$, the highest among all models. This indicates that HiPPrO is not only attentive to the target group but also maintains a strong semantic connection to the original content, making its CS both specific and contextually relevant. Furthermore, when examining the toxicity scores, HiPPrO achieves a respectable score of $0.087$, which, although slightly higher than GPT-3.5-Turbo Zero-Shot, still indicates a significantly low presence of harmful language comparable to all other baselines. This suggests that HiPPrO effectively balances generating contextually rich CS while keeping the content non-toxic and constructive. Statistical analyses reveal that our model significantly outperforms both GPT-3.5 ZS and FS across most metrics, with exceptions in meteor, CoSim, and TC scores. Please refer to Appendix (Section 8.5) for additional information about statistical significance tests.
6.2 Ablation Study
Our ablation study assesses the effects of several components of HiPPrO (Table 2). One such study involves fine-tuning HiPPrO without the ORPO component. The results show that the omission of ORPO leads to a decrease in performance in almost all evaluation metrics, with a notable reduction in the toxicity score by $0.024$. This suggests that ORPO contributes significantly to enhancing both the overall performance and the generation of less toxic counterspeech. Additionally, we fine-tune HiPPrO with DPO instead of ORPO and observe a slight degradation in performance. Although the difference is marginal, the key advantage of ORPO over DPO lies in ORPO’s ability to operate without a reference model, which DPO requires. Additionally, we investigate how different numbers of VT in HiPPrO impact its performance, testing configurations with VT= $5,7,$ and $10$. The results indicate that there is no notable performance gain after VT=$5$. While higher virtual tokens lead to a slight reduction in the toxicity score, the improvements are insufficient across other metrics, suggesting no significant returns with increased token counts.
Our ablation studies also include evaluations of Vanilla FLAN-T5XXL and Vanilla BARTLarge, both with prefix tuning, without hierarchical learning, providing further insights into HiPPrO’s effectiveness. We observe that while these models offer competitive performance, they do not surpass HiPPrO in several key metrics. For Vanilla FLAN-T5XXL, prefix tuning yields ROUGE-1, ROUGE-2, and ROUGE-L scores of $0.229$, $0.052$, and $0.158$, respectively. These scores are slightly lower than those of HiPPrO, indicating that while the prefix-tuned FLAN-T5XXL performs well, HiPPrO’s output is more detailed and comprehensive. However, the IC scores for Vanilla FLAN-T5XXL and BARTLarge are $0.470$ and $0.249$, respectively, while their EC scores are $0.666$ and $0.453$, respectively. HiPPrO achieves higher IC and EC scores ($0.929$ and $0.706$ respectively), highlighting the limitations of simple prefix-tuning methods in effectively aligning CS with multiple attributes.
6.3 Human Evaluation
Previous studies (Jones et al., 2024; Wang et al., 2023) emphasize the need for a dual evaluation framework, as automatic metrics show weak correlation with human judgments of counterspeech effectiveness. A comprehensive human evaluation was conducted on a random subset of responses from the top-performing CS generation methods (HiPPrO, CoARL, Few-Shot GPT-4, and GPT-3.5 Turbo), ensuring uniform distribution across intents and emotions. A diverse panel of 35 experts in NLP and social sciences (aged 20-35, 45% male, 55% female) evaluated and ranked these responses based on several key metrics. We followed Hengle et al. (2024) for our human evaluation. The evaluation framework consists of five key metrics: Independent Counterspeech (ICS) to gauge the response’s self-sufficiency; Adequacy (Ad) to assess its linguistic quality; Contextual Relevance (CoRl) to measure its responsiveness to hate speech components; and Argumentative Effectiveness (ArgE) to evaluate its carefulness and convincing. We present the comparative performance of HiPPrO against leading methods through Win Rate scores (see Table 3). For ICS, HiPPrO outperforms the baselines with win rates of $0.91$ and $0.89$ over GPT-4 and GPT-3.5, showing its ability to generate CS that operates effectively without needing extra context. Regarding Ad, HiPPrO achieves higher scores of $0.85$ and $0.87$, reflecting the superior grammatical accuracy and fluency of CS. In terms of CoRl, HiPPrO’s win rates of $0.89$ and $0.87$ highlight its strength in addressing crucial aspects of hate speech, such as targeted biases. Finally, for ArgE, HiPPrO leads with scores of $0.89$ and $0.90$, indicating its effectiveness in delivering compelling and well-structured CS. These results collectively show HiPPrO’s robust performance across all metrics compared to the baseline models.
Models on comparison Metrics
ICS $\uparrow$ Ad $\uparrow$ CoRl $\uparrow$ ArgE $\uparrow$
HiPPrO vs CoARL $0.96$ $0.98$ $0.93$ $0.97$
HiPPrO vs GPT-4 (FS) $0.91$ $0.85$ $0.89$ $0.89$
HiPPrO vs GPT-3.5 (FS) $0.89$ $0.87$ $0.87$ $0.90$
Table 3: Results of the human evaluation study, where responses generated by HiPPrO are shown against those produced by (a) CoARL, (b) GPT-4 (FS), and (c) GPT-3.5 (FS). The results are reported in terms of Win Rate %, indicating the % of instances where HiPPrO outperforms the respective baselines.
7 Conclusion
This study presented HiPPrO, a novel two-stage framework for generating controllable, multi-attributed counterspeech. The initial stage comprised a hierarchical learning process, where the model acquired attribute-specific prefixes, thereby guiding the LLM towards targeted counterspeech generation. The subsequent stage involved refining the outputs to enhance their human-like quality and non-toxicity, employing a reward- and reference-free alignment approach. Additionally, we introduced the multiCONAN dataset with intent- and emotion-specific counterspeech. An extensive evaluation, incorporating a range of quantitative and qualitative measures, demonstrated HiPPrO’s superiority over multiple baselines.
Limitation
Our research presents several limitations that warrant consideration. Firstly, the dataset utilized for hate speech and counterspeech is not comprehensive, potentially omitting various forms and targets of online hate. Secondly, the framework’s reliance on pre-trained models may introduce inherent biases or inaccuracies stemming from these source models. Additionally, the evaluation metrics employed do not fully align with human perceptions of counterspeech quality, thereby failing to capture the intricate nuances of natural language interactions. Moreover, our framework does not address the possibility of feedback loops or escalation that could arise following the generation of counterspeech, which may influence the long-term effectiveness and impact of our approach. Lastly, while efforts were made to maintain high-quality annotations for counterspeech, it is conceivable that our dataset may not match the caliber of those annotated by more experienced operators from NGOs, such as those found in the Multi-Target CONAN Fanton et al. (2021b) dataset. Future research could mitigate these limitations by expanding and diversifying the dataset, enhancing the reward function and evaluation criteria, and incorporating dialogue modeling into the framework.
Ethics Statement
We recognize the sensitivity required in addressing online hate speech and acknowledge the ethical and moral complexities inherent in conducting research in this area. This initiative serves as an initial attempt to compile a comprehensive and varied collection of counterspeech responses for each instance of hate speech encountered. We understand that algorithms developed for automated counterspeech may generate responses that fail to accurately convey the intended meanings, highlighting the urgent need to better integrate real-world knowledge into these systems. Despite the potential of generative algorithms, there remains a critical necessity for a robust and diverse database of counterspeech to ensure consistently favorable outcomes. Furthermore, while fully operational counterspeech algorithms have yet to be realized, organizations such as United Against Hate play a crucial role in mitigating the prevalence of hate speech in online environments.